a) 2 b) 5   c) 3   d)1 e)4

RnADzFvsJkEAAAAASUVORK5CYII=

a) 2 b) 5   c) 3   d)1 e)4
Answer #1

total energy of oscillation = k A^2 /2

= 20(2^2)/2

= 40 J


friction force = u m g = 0.2 x 1 x 9.8 = 1.96 N

energy dissipated till first pass = 1.96 x 2 = 3.92 J

first to second pass = 1.96 x 4 = 7.84 J


total pass = 1 + ((40-3.92)/7.84) = 5.6

An(B) 5 pass